Miguel is now twice as old as Pauline. Five years ago Miguel was three times as old as Pauline. Find their present ages.

Answer by ikleyn(52798)


Let  x  be the Pauline's present age.

Then the Miguel's present age is  2x,  in accordance with the condition.

Five years ago Miguel's age was  (2x-5)  and Pauline's age was  (x-5).

In accordance with the condition,  you have an equation

2x-5 = 3*(x-5).

Solve it step by step:

2x - 5 = 3x - 15,
-5 + 15 = 3x - 2x,
10 = x.

Thus Pauline's present age is  10 years.
Hence,  Miguel's present age is  2*10 = 20 years.